OpenStudy (anonymous):

a kirby vacuum cleaner dealer ship sold 4300 units in 2006 and 407 units in 2007. find the percent increase or decrease in the number of units sold.

hero (hero):

Hint: 4300 + 470 = 8370 total units sold 2nd Hint: \[\% \Delta = \frac{|a - b|}{a} \times 100\] a = first number b = second number